ect.mheducation.com/flow/connect.html?returnUrl=https%3A%2F%2Fconnect.mheduca Oblem 9-24 i participants to avoid some or allCheck my work mode: This shows what is correct or incorrect for the work you have completed so far. It does not or favorableExchange Corp. is a company that acts as a facilitator in tax-favored real estate swaps. Such swaps, know as 1031 exchanges, permit participants to avoid some or all of the capital gains taxes that would otherwise be due. The bookkeeper for the company has been asked to prepare a report for the company to help its owner/manager analyze performance. The first such report appears below: Exchange Corp. Analysis of Revenues and Costs For the Month Ended May 31 Actual Unit Revenues and Costs Planning Budget Unit Revenues and Costs Variances Exchanges completed 30 25 Revenue $ 620 $ 720 $ 100 U Expenses: Legal and search fees 263 240 23 U Office expenses 227 260 33 F Equipment depreciation 25 30 5 F Rent 75 90 15 F Insurance 15 18 3 F Total expense 605 638 33 F Net operating income $ 15 $ 82 $ 67 U Note that the revenues and costs in the above report are unit revenues and costs. For example, the average office expense is $260 per exchange completed on the planning budget; whereas, the average actual office expense is $227 per exchange completed. Legal and search fees is a variable cost; office expenses is a mixed cost; and equipment depreciation, rent, and insurance are fixed costs. In the planning budget, the fixed component of office expenses was $5,100. All of the company’s revenues come from fees collected when an exchange is completed. Required: 1. Is the report prepared by the bookkeeper useful as a performance report? 2. Complete the performance report that would help the owner/manager assess the performance of the company in May.
Answer #1
1
NO, the report prepared by bookkeeper is not a useful performance report
2
Actual results Revenue and spending variances Flexible budget Activity variances Planning budget
Exchanges completed 30 30 25
Revenue 18600 3000 U 21600 3600 F 18000
Expenses:
Legal and search fees 7890 690 U 7200 1200 U 6000
Office expenses 6810 30 U 6780 280 U 6500
Equipment Depreciation 750 0 None 750 0 None 750
Rent 2250 0 None 2250 0 None 2250
Insurance 450 0 None 450 0 None 450
Total expense 18150 720 U 17430 1480 U 15950
Net operating income 450 3720 U 4170 2120 F 2050
Office expenses:
Variable component = (6500-5100)/25 = $56
Office expenses in flexible budget = 5100+(30*56) = $6780